# EasyText: Controllable Diffusion Transformer for Multilingual Text Rendering

### v1

@@ 3.5 EasyText Dataset Construction
High-Quality Human-Annotated Dataset. The second dataset comprises 20k high-quality image–text box pairs, primarily in Chinese and English. Text regions are extracted using PP-OCR [10] and filtered to ensure annotation quality. In these examples, text is more naturally integrated into scene content, often with stylized typography and visually consistent stroke patterns. The rendered text is not included explicitly in the prompt, but instead triggered by placeholder tokens (e.g., <sks1>, <sks2>), supporting semantic alignment between text and background. This dataset facilitates better visual-text fusion and improves typographic aesthetics.
4 Experiments
4.1 Experimental Settings
Evaluation Metric. To evaluate the accuracy of multilingual text generation, we adopt two complementary metrics: (i) Character-level precision which measures the correctness of generated characters relative to ground-truth texts. This reflects the model’s ability to produce accurate predictions at the most basic unit level for each language. (ii) Sentence-level precision, which assesses whether the entire text line content within each controlled text box is rendered completely correctly. This captures the overall consistency and correctness at the region level.
Multilingual Benchmark. We construct a multilingual benchmark with 90 language-agnostic prompts across 10 languages. For each language, prompts are paired with language-specific text to generate condition images while preserving semantic intent. Four images are generated per prompt-language pair. For logographic scripts (Chinese, Japanese, Korean), the average number of characters per text box is 7.1, and for alphabetic scripts, it is 14.3. Each image contains an average of 1.7 text boxes. To evaluate the generated results, three volunteers were tasked with assessing the accuracy of the text in the images, ensuring reliable multilingual generation quality.
@@ 4.1 Experimental Settings
Korean 4308 13 100K 0.1K
Japanese 2922 17 100K 0.1K
Thai 2128 3 100K
Vietnamese 91 19 100K
Latin 104 30 180K 16K
Greek 79 19 100K
@@
Text-Image Fusing 74.48 80.28 74.80 81.12 78.66
4.2 Comparision results
Multi-lingual Precision Evaluation. We evaluate our model on a multilingual benchmark against state-of-the-art commercial models (e.g., FLUX [20], Jimeng AI [5], SD3.5 [11]) and other methods (e.g., Glyph-ByT5-v2 [25], AnyText [45]), following the benchmark and metrics outlined in Sec. 4.1. For languages with limited support in prior rendering methods, we compare primarily with Glyph-SDXL-v2. All baseline models are evaluated using their official inference settings. For each sample, four images are generated with identical image descriptions and target texts across all methods. For models lacking conditional input support, the target text is included in the prompt; Evaluation focuses solely on text accuracy, regardless of spatial layout or surrounding content.
Figure 5: Comparison of text rendering on the pretraining synthetic dataset using multiple font overlays versus single font overlays. The red boxes highlight the erroneous regions.
The results show that our model surpasses competing methods in text rendering precision across several languages, including English and Italian (Table 5) and achieves strong character-level accuracy in Chinese, though slightly behind Jimeng AI at the sentence level. It also performs well in unsupported languages like Thai and Greek. In Table 5, it also leads in OCR accuracy and improves over FLUX in CLIPScore, indicating higher visual-text alignment. These trends are further supported by GPT-4o evaluation results. After pretraining, the model demonstrates strong text rendering performance from condition images. However, it shows limited text-image coherence, as indicated by lower CLIPScore and GPT-4o evaluations. Fine-tuning alleviates this issue, significantly improving CLIPScore, Text Aesthetics, and overall visual-text alignment. Although rendering accuracy slightly declines due to flexible positioning and artistic font usage, the trade-off is acceptable given the qualitative gains. Notably, while most competing methods use datasets exceeding 1M samples, our fine-tuning relies on only  20k annotated examples. Moreover, the large-scale pretraining set can be replaced with synthetic data, reducing dependence on large, annotated datasets.
4.3 Qualitative Results
We provide qualitative comparisons of multilingual text rendering, including Chinese, English, and other languages. Results are shown in Fig. 4. Compared to previous region-controlled text rendering methods, our approach demonstrates significant improvements in visual quality and text fidelity, with better visual-text integration, higher OCR accuracy, and enhanced aesthetic coherence (Shown in Fig. 3). Additionally, unlike commercial text-to-image models, our method allows for more precise spatial control over rendered text, particularly in generating multiple paragraphs of relatively long text (e.g., 4–5 paragraphs with around 20 characters each) while maintaining consistent layouts. We further evaluate the model’s ability to handle challenging text generation scenarios that are typically difficult for existing methods. Specifically, our approach generalizes well to unseen characters, slanted or non-rectilinear text regions, and long text sequences, while maintaining structural consistency and legibility. Detailed qualitative results are provided in the appendix.
Figure 6: Ablation study comparing the full conditioned method with: (1) layout-free (2) without condition inputs, and (3) original FLUX. The red boxes highlight the erroneous regions.
@@ 4.4 Ablation Studies
To validate the effectiveness of using diverse fonts in synthetic pretraining, we conducted a controlled experiment where both the target and condition images were rendered using the same standard font. While this setup yields high initial precision, it mainly captures a direct shape mapping and fails to generalize after fine-tuning on real-world data(Shown in Fig. 5). This highlights the importance of font diversity—rendering targets with multiple fonts while keeping the condition font fixed—for learning robust, transferable representations. we compare our full approach—using both the condition input and implicit character position alignment-with three ablated variants: (1) removing position alignment, (2) removing the condition input, and (3) using the original FLUX model. As shown in Fig. 6, incorporating position mapping substantially improves spatial precision and text accuracy, particularly in multi-text scenarios. In comparison, when removing the condition input, the target render text is instead provided in the prompt(in our full method, renter text content is given via the condition input). This demonstrates that our fine-tuned model effectively retains the strong generation capability of the base FLUX model.
@@ 5 Limitation
The Implicit Character Position Alignment mechanism is less effective when character positions are substantially overlapping, occasionally leading to reduced rendering accuracy. In addition, training across multiple scripts results in confusions between simple but visually similar characters from different writing systems, such as rendering the Vietnamese character "" as "e". These cases are infrequent but consistently observed.
@@ 6 Conclusion
In this work, we propose EasyText, a diffusion-based framework for multilingual and controllable text generation in text-to-image synthesis. The method learns to mimic font rendering by mapping glyph features into the same latent space as the image via a VAE, and achieves Implicit Character Position Alignment by combining position encoding with spatial targets, enabling precise and flexible text placement. Benefiting from a two-stage training strategy—pretraining on a large-scale synthetic multi-font dataset followed by fine-tuning—EasyText reduces the need for real multilingual data while maintaining high rendering accuracy. It also demonstrates strong visual-text integration, effectively embedding text into complex scenes. Extensive experiments confirm the effectiveness and generalization ability of our approach.
Supplementary Material for EasyText: Controllable Diffusion Transformer for Multilingual Text Rendering
David S. HippocampusUse footnote for providing further information about author (webpage, alternative address)—not for acknowledging funding agencies.
Department of Computer Science
Cranberry-Lemon University
Pittsburgh, PA 15213
hippo@cs.cranberry-lemon.edu
Appendix
This supplementary material is organized into three sections. Section A describes the training datasets used for pretraining and fine-tuning, including illustrative examples. Section 12 presents diverse and comprehensive generated examples. Section C outlines the process and sample cases of the GPT-based evaluation and user study.
Appendix A Additional Dataset Details
The main paper provides a detailed description of the dataset’s design and composition. This section complements that discussion by offering additional details and presenting example instances from the dataset to further illustrate its structure.
Figure 7: Sample images from the synthetic dataset used in the pretraining.
A.1 Synthetic Dataset for Large-Scale Pretraining
For the training dataset used in pretraining, we collected background images from several publicly available datasets without textual content, such as LAION-Aesthetics. For each script (i.e., a language-specific character set), we rendered text in various fonts onto these backgrounds. During the rendering process, we recorded both the text content and the corresponding bounding box coordinates, producing synthetic image–text pairs designed for text rendering tasks. This approach enabled the creation of a large-scale multilingual dataset of approximately 1 million samples, offering a practical alternative for low-resource languages, where high-quality annotated image–text pairs are often unavailable. Some sample instances are shown in Fig. 7.
A.2 High-Quality Dataset for Fine-Tuning
For the fine-tuning stage, we collected around 20K images containing text from copyright-free websites. Text regions were detected and filtered using PP-OCR, and each image was further annotated with a caption generated by GPT-4 Mini. Some sample instances are shown in Fig. 8.
Figure 8: Sample images from the high-quality dataset used in the fine-tuning.
Appendix B Additional Generation Results
We present additional diverse generation results produced by EasyText (Fig. 9). Fig. 10 and Fig. 11 further illustrate the model’s ability to generate text in multiple languages, all achieved using a single, general-purpose model. Additional figures also highlight other powerful capabilities of EasyText (Fig. 12), including layout-free text generation, adaptation to curved or slanted regions, and strong generalization to unseen characters and languages—both of which were never encountered during training.
Figure 9: Additional text rendering results generated by EasyText
Figure 10: Results of EasyText across multiple languages, showing the first group of languages.
Figure 11: Results of EasyText across multiple languages, showing the second group of languages.
Figure 12: Additional results generated by EasyText, showcasing unique capabilities beyond existing methods, such as handling unseen characters and languages.
Appendix C Additional Evaluation Details
We selected several metrics to assess the quality of text rendering results, including Image Aesthetics, Text Aesthetics, Text Quality, and Text-Image Fusion. These metrics were evaluated through two distinct evaluation methods: GPT-based automatic evaluation and a user study.
C.1 GPT-Based Evaluation
Figure 13: Example dialogue illustrating the evaluation process using GPT-4o.
The evaluation procedure for an individual image using GPT is illustrated in Fig. 13. The figure shows an example of the dialogue-based evaluation process for clarity; in practice, the evaluations were conducted by directly querying the GPT-4o API, which automatically returned the scores. Evaluations were performed at scale, and the final scores for each metric were computed as the average across all samples.
C.2 User Study Design
Table 6: User study results (%): Comparison of different methods based on user preferences across four evaluation metrics. The percentages represent the average proportion of times each method was selected as the best-performing one across multiple groups of samples.
Metrics EasyText (pretrain) EasyText (fine-tune) Glyph-v2 SD3.5 FLUX AnyText
Image Aesthetics 15.05 25.01 16.67 21.31 7.58 14.39
Font Aesthetics 15.91 31.06 9.85 15.15 13.64 14.39
Text Quality 11.65 22.97 14.39 21.10 13.96 15.91
Text-Image Integration 15.15 34.09 14.62 11.36 12.64 12.14
The user study followed a questionnaire-based format, as detailed in Fig. 14, which illustrates an example group of comparison images presented in the questionnaire. Approximately 50 questionnaires were distributed. Each questionnaire compared outputs from different models across ten groups of generated results.The results of the user study are shown in Table 6.
Figure 14: An example group of comparison images from the user study questionnaire.

### v2

@@ 3.4 EasyText Dataset Construction
Korean 4308 13 120K 0.1K
Japanese 2922 17 120K 0.1K
Thai 2128 3 120K
Vietnamese 91 19 120K
Latin 104 30 180K 15K
Greek 79 19 120K
High-Quality Human-Annotated Dataset. The second dataset consists of approximately 20k high-quality image–text box pairs, primarily in Chinese and English. Text regions are extracted using PP-OCR (Du et al. 2020) and filtered to ensure annotation quality. In these examples, text is more naturally integrated into scene content, often with stylized typography and visually consistent stroke patterns. The rendered text is not included explicitly in the prompt, but instead triggered by placeholder tokens (e.g., <sks1>, <sks2>), supporting semantic alignment between text and background. This dataset facilitates better visual-text fusion and improves typographic aesthetics.
4 Experiments
4.1 Experimental Settings
Evaluation Metric. To evaluate the accuracy of multilingual text generation, we adopt two complementary metrics: (i) Character-level precision, which measures the correctness of generated characters relative to ground-truth texts. This reflects the model’s ability to produce accurate predictions at the most basic unit level for each language. (ii) Sentence-level precision, which assesses whether the entire text line content within each controlled text box is rendered completely correctly. This captures the overall consistency and correctness at the region level.
Multilingual Benchmark. We construct a multilingual benchmark with 90 language-agnostic prompts. For each language, prompts are paired with language-specific text to generate condition images while preserving semantic intent. Four images are generated per prompt-language pair. For logographic scripts, the average number of characters per text box is 7.1, and for alphabetic scripts, it is 14.3. Each image contains an average of 1.7 text boxes. To evaluate the generated results, three volunteers were tasked with assessing the accuracy of the text in the images, ensuring reliable multilingual generation quality.
4.2 Comparison Results
@@
We evaluate our model on a multilingual benchmark against state-of-the-art commercial models (e.g., FLUX (Labs 2024), Jimeng AI (ByteDance 2024), SD3.5 (Esser et al. 2024)) and other methods (e.g., Glyph-ByT5-v2 (Liu et al. 2024b), AnyText (Tuo et al. 2023)), following the benchmark and metrics outlined in Sec. 4.1. For languages with limited support in prior rendering methods, we compare primarily with Glyph-SDXL-v2. All baseline models are evaluated using their official inference settings. For each sample, four images are generated with identical image descriptions and target texts across all methods. For models lacking conditional input support, the target text is included in the prompt; evaluation focuses solely on text accuracy, regardless of spatial layout or surrounding content.
The results show that our model surpasses competing methods in text rendering precision across several languages, including English and Italian (Table 4) and achieves strong character-level accuracy in Chinese (Table 3), though slightly behind Jimeng AI at the sentence level. It also performs well in unsupported languages like Thai and Greek. In Table 5, it also leads in OCR accuracy and improves over FLUX in CLIPScore, indicating higher visual-text alignment. These trends are further supported by GPT-4o evaluation results. After pretraining, the model demonstrates strong text rendering performance from condition images. However, it shows limited text-image coherence, as indicated by lower CLIPScore and GPT-4o evaluations. Fine-tuning alleviates this issue, significantly improving CLIPScore, Text Aesthetics, and overall visual-text alignment.
@@
4.3 Qualitative Results
We provide qualitative comparisons of multilingual text rendering, including Chinese, English, and other languages. Results are shown in Fig. 4. Compared to previous region-controlled text rendering methods, our approach demonstrates significant improvements in visual quality and text fidelity, with better visual-text integration, higher OCR accuracy, and enhanced aesthetic coherence (shown in Fig. 3). Additionally, unlike commercial text-to-image models, our method allows for more precise spatial control over rendered text, particularly in generating multiple paragraphs of relatively long text (e.g., 4–5 paragraphs with around 20 characters each) while maintaining consistent layouts. We further evaluate the model’s ability to handle challenging text generation scenarios that are typically difficult for existing methods. Specifically, our approach generalizes well to unseen characters, slanted or non-rectilinear text regions, and long text sequences, while maintaining structural consistency and legibility. (refer to Appendix Section C for details)
@@ 4.3 Qualitative Results
Text-Image Fusing 74.48 81.28 74.80 81.12 78.66
@@ 4.4 Ablation Studies
Figure 5: Ablation study comparing the full conditioned method with: (1) layout-free (2) without condition inputs, and (3) original FLUX-Dev. <sks1>, <sks2> triggers are used to replace rendered text in the first two cases.
To validate the effectiveness of using diverse fonts in synthetic pretraining, we conducted a controlled experiment where both the target and condition images were rendered using the same standard font. Despite high initial precision, this setup captures only direct shape mappings and fails to generalize after real-world fine-tuning. (refer to Appendix Section E) This highlights the importance of font diversity—rendering targets with multiple fonts while keeping the condition font fixed—for learning robust, transferable representations. We compare our approach with three ablated variants: (1) removing position alignment, (2) removing the condition input, and (3) using the original FLUX model. As shown in Fig. 5, incorporating position mapping substantially improves spatial precision and text accuracy, particularly in multi-text scenarios. In comparison, when removing the condition input, the target rendered text is instead provided in the prompt(in our full method, rendered text content is given via the condition input). This demonstrates that our fine-tuned model effectively retains the strong generation capability of the base FLUX model.
@@ 5 Limitation
The Implicit Character Position Alignment mechanism is less effective when character positions are substantially overlapping, occasionally leading to reduced rendering accuracy. In addition, training across multiple scripts results in confusion between simple but visually similar characters from different writing systems. These cases are infrequent but consistently observed.
@@ 6 Conclusion
In this work, we propose EasyText, a diffusion-based framework for multilingual and controllable text generation in text-to-image synthesis. EasyText learns to mimic glyph features and supports precise and flexible text placement through implicit character alignment. With a two-stage training strategy, the method significantly reduces reliance on real multilingual data while maintaining high rendering accuracy. It also demonstrates strong visual-text integration, effectively embedding text into complex scenes.